The number of fish that jon catches is dependent upon the lures he uses. He also uses lots of expensive lures. Here in S.C each fisherman can use 2 poles at one time so jon can use two different lures. If L^1 represents the number of fish caught per hour for the first lure on the pole that he holds and if L^2 represents the number of fish caught per hour for the lure on the pole that he puts on the "dummy" holder, then the formula below represents the relation between the two lures used and the total number of lures lost per hour using the 2 poles (C is the rate of lures lost per hour using both poles) If the combined lost rate and the rate for fish on lure 2 are known, write a function for the rate of fish on the first lure (solve the equation for L^1)
1/C=1/L^1+L^2/L^1

Answer by stanbon(75887)


The number of fish that jon catches is dependent upon the lures he uses. He also uses lots of expensive lures. Here in S.C each fisherman can use 2 poles at one time so jon can use two different lures. If L1 represents the number of fish caught per hour for the first lure on the pole that he holds and if L2 represents the number of fish caught per hour for the lure on the pole that he puts on the "dummy" holder, then the formula below represents the relation between the two lures used and the total number of lures lost per hour using the 2 poles (C is the rate of lures lost per hour using both poles) If the combined lost rate and the rate for fish on lure 2 are known, write a function for the rate of fish on the first lure (solve the equation for L1)
1/C=1/L1+L2/L
Multiply thru by L1 to get:
L1/C=1+L2
Multiply thru by C to get:
L1=C(1+L2)
or
L1=C+C(L2)
---------------
Cheers,
Stan H.